wiki:UC20201110

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 11/10/2020

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Denise Stephens (BYU), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Sarah Tuttle (UW), Chip Kobulnicky (UWy), Derek Buzasi (FGCU), Jeff Bary (NAPG)


User feedback and comments from institutional representatives

  • Georgia State - nothing to report
  • Seattle - nothing to report
  • Washington - nothing to report
  • Oklahoma - user reported a tracking issue in October and wondered about the status; Jamey said that Bill fixed some software/driver issues that he thinks addressed this. Please let us know if these issues continue to be a problem.
  • NAPG - nothing to report
  • JHU - nothing to report
  • CU - nothing to report
  • UVa - nothing to report
  • NMSU - nothing to report
  • BYU - nothing to report
  • FGCU - nothing to report
  • Wyoming - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 10/01/2020 - 11/05/2020

0) Overview

APO has been open for night operations since June 1. We are still restricting all visitors to the site. Following the state of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors.

Currently, as NMSU is in Phase 3 of the Return to Research, APO is allowed between 35-50% of staffing. The NMSU numbers, coupled with the state of New Mexico restrictions for a non-essential classified business like APO, which allow staffing levels of 50% of the Fire Marshall determined maximum occupancy of a building, has allowed us to have more people on site at any given time. However, those restrictions are still limiting work activities for daytime engineering tasks.

There has been a dramatic increase of COVID-19 cases in New Mexico in the last 4-6 weeks (from ~90 cases per day for the 7-day average in mid-September to 915 cases per day for the 7-day average currently). Combined with the dramatic increase in New Mexico, there has been a similar increase in El Paso county, which has led to the restriction that NMSU employees who live in El Paso county can’t be asked to travel into New Mexico to work.  Also, El Paso County is currently under a 2 week stay at home order that runs through 11/12.

Weather for the month of October has been as expected for an average year. The Miller moth counts are decreasing considerably, associated with the first cold weather/snow that occurred the week before Halloween.

1) Telescope

The telescope has been performing as expected. An issue with the altitude roller bearings, which had caused occasional tracking problems and increased wind sensitivity through the summer, was diagnosed after the early October engineering time and parts were replaced. Further tests in late October and early November showed distinct improvements in pointing, tracking, and robustness against wind.

2) Instruments

DIS: Regarding the ongoing scattered light issue, the red camera is stable, and the blue camera returned from servicing and is holding relatively stable at a usable level.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The inter-order light (IOL) for red and blue continues to improve slowly with cooling weather.

NIC-FPS: Frequent expose inits are being required to remedy an all-zero readout problem. Investigations are ongoing.

ARCTIC: powered on, cold, and operational. An issue with dome flats is being debugged and solutions are being worked on.  Specifically, a “fence” baffle on the mirror cover petal design is being worked on; initial design ideas had to be modified because of interference issues with the telescope structure. The new rotator mechanism for the diffuser arrived on site and installation will begin after November 17.

3) ARCSAT

ARCSAT is operational and has been used for observations this past month. Driver/software issues associated with changing temperature as well as upgraded software have settled down and appear to be resolved.   The Shelyak eShel commissioning took place at the beginning of September, the end of October and the first week of November.  The data from these commissioning runs are currently being analyzed and discussed.


Additional telescope and instrument discussion

Not many updates regarding telescope or instruments beyond what was included in the monthly report. The site is still operating at a reduced on-site staffing level due to COVID restrictions (discussed further below). Some dome roller bearings were replaced. DIS Blue shows less scattered light than in the recent past so it is currently performing better than expected. The ARCES IOL is improving. We are still investigating why NIC-FPS requires expose-inits. The "fence baffle" is currently being worked on to mitigate scattered light in ARCTIC dome flats. The tracking issue with ARCSAT is hopefully resolved.


APO and the COVID-19 virus

The COVID situation in southern New Mexico and El Paso is quite severe. El Paso is under a stay-at-home order, and hospitals in New Mexico are nearly full. Both areas have high positivity rates. Users should be aware that there is a non-zero chance that we will need to shut down again, as we are already operating short-handed due to quarantining/travel restrictions for some of the staff. Visitors are still prohibited from coming to the site. Although it is unclear how long these restrictions will remain in place, groups planning class training or observing trips should expect them to be conducted remotely through Q1 2021. [Note added in proof: the observatory was shut down between Nov 16 - Dec 01 per a public health order from the Governor of New Mexico. The 3.5m resumed science operations on Dec 2 and the observatory is currently open, albeit at a 25% on-site staffing level.]


Q4 3.5m scheduling

The Q4 schedule is posted. There is some OPEN time available, particularly for B-half on Thanksgiving, Nov 29A, and some time in mid/late December. UC reps: notify users that there is a limited amount of OPEN time Q4. If interested in requesting some time, please follow standard protocols for making such requests.


Q4 ARCSAT scheduling

The Q4 schedule is full. A call for Q1 ARCSAT proposals will go out after the Q1 3.5m schedule is complete, so probably in early/mid-December.


Q1 3.5m scheduling

The Q1 2021 allocations were sent out last week, and the proposals are due on 11/25. The dates for the Q1 training session are still TBD. March can sometimes be heavily subscribed, so institutional schedulers are asked to elevate the priority of those programs who MUST be scheduled in March.


Updated Training Policies and Procedures

As a reminder, we are now offering one training session per quarter for PhD-level users who need to be trained for remote observing (postdocs, research staff, new faculty, etc.). The first pickup training session, held in Q4, was well attended, and we learned some lessons for how to improve the trainings for future groups. Note that these sessions are not intended for students except in extenuating circumstances. Students should work directly with their advisors (who should propose for observing time through the normal channels) to get trained; advisors should let Ben, Russet, and Nancy know if there are any students in their observing teams who require training.


KOSMOS Update

This week several members of the TEG are getting trained on FARO arm, which will enable more precise optical alignments that will be required for the installation of the CCD into the dewar. The remote installation teams working on the 2.5m seem to be having success, so the TEG is pushing for more KOSMOS commissioning work to be done in lab rather than on the telescope. The working assumption is that most (or perhaps all) of the TEG members are not likely to come to APO for KOSMOS commissioning given the current COVID-19 situation, so they are developing plans for how to do things remotely. They are developing inputs for the pre-ship review so that is clear what the expectations are for tasks that can be done in lab versus on sky.


ARCTIC Diffuser Rotation

The diffuser rotation mechanism is scheduled for commissioning to begin on November 17.


Science Communications Update

There is an APOLLO outreach event scheduled on November 19; this will be live-streamed through the Las Cruces Museums Facebook page.


ARC Board of Governors Meeting

The ARC Board voted to approve the 3.5m budget. The main input from the Board was that they would like to see the establishment of an instrumentation working group that can assess the needs of the consortium members concerning an echelle replacement and provide recommendations for potential pathways for doing that. Nancy will be constituting that group in the near future.


Planning for Spring Semester

After confirming that there were no severe schedule conflicts, it was agreed upon that User Committee meetings will remain on the first Tuesday of the month at 10:30 am Mountain Time for the spring semester. Nancy would like to move forward with planning a one-day science meeting highlighting results using APO. This will be a virtual meeting; it was recommended by several UC members that this be scheduled during the summer.


ACTION ITEMS

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. She is trying to reconcile some vastly different price quotes from several vendors and is looking to place the order within the next month.
  • UC reps: remind your users that if they need to download their data there is a read-only FTP site from which they can do so without needing the institutional password for the current quarter. Status: CLOSED. Presumably this has been communicated and this will be removed from the action items list.
  • UC reps: notify users that there is a limited amount of 3.5m OPEN time in Q4. If interested in requesting some time, please follow standard protocols for making such requests.
  • Nancy: develop a survey to get feedback on the pickup training session. Status: OPEN.

New action items from this meeting:

None.


All Other Business

None.


Next meeting

The next meeting will be on Tuesday December 8 at 10:30 am MST.


Note: See TracWiki for help on using the wiki.